# Dynamic Matroids: Base Packing and Covering

### v1

@@ A.2 Base Covering
Let $E_{i}:=E\setminus\bigcup_{j<i}A_{j}$ and let $\mathcal{M}_{i}=\mathcal{M}|E_{i}$ for all recursion levels denoted by the indices $i\in I$. First, we note that $1/\min_{e}\ell^{*}(e)=\max_{i\in I}\Phi_{\mathcal{M}_{i}}$ over all recursion levels $i$. Let $j$ be the last iteration in which $\Phi_{\mathcal{M}_{i}}$ is maximized. Since the $\Phi$-values are non-decreasing, see Lemma˜A.2, this also has to be the last level of the recursion and therefore $E_{j+1}=E_{j}\setminus A_{j}=\emptyset$ and $\operatorname{rk}(E_{j})=\operatorname{rk}(A_{j})$. We get
@@
$|Y|=\sum_{i\in I}|A_{i}\cap Y|\leq\sum_{i\in I}\Phi_{\mathcal{M}_{i}}\operatorname{rk}(A_{i}\cap Y)$
where it remains to show the inequality $|A_{i}\cap Y|\leq\Phi_{\mathcal{M}_{i}}\operatorname{rk}(A_{i}\cap Y)$ for any $i\in I$. Assume towards a contradiction that there is a level $k\in I$ for which the inequality does not hold. In the following let $\overline{A_{i}}:=E_{i}\setminus A_{i}$ for each $i\in I$. Now, consider the value
@@
since the matroid $\mathcal{M}_{i}$ contains $\Phi_{\mathcal{M}_{i}}$ disjoint bases. Hence, every base contains at most $|A_{i}\cap Y|/\Phi_{\mathcal{M}_{i}}$ many elements. Next, we want to show
@@
Appendix B Greedy Base Collection
First, let us define a greedy base collection as follows: let the weight of an element be the number of bases an element belongs to, and require that the base in the collection form successive minimum weight bases. This is a natural generalization of the notion for graphic matroids called ‘greedy tree-packing’ [^5] We opt for ‘collection’ instead of ‘packing’ to avoid any confusion with base packing as in Theorem A.1., see, e.g., [Thorup07].
@@ Appendix B Greedy Base Collection
B.1 Dynamic Greedy Base Collection
@@
Given a greedy base collection $\mathcal{B}$, let there be an insertion or deletion $e$. The element $e$ was/will be part of at most $|\mathcal{B}|$ bases. For each base it was/will be a part of, we get at most $|\mathcal{B}|$ changes to other bases. Note that this is non-trivial; at a first glance, the changes over bases could add up. However, we note that if an element $f$ gets an increased weight because it was used in some base, and later this increased weight means that in another base $g$ was chosen instead of $f$, then the weight of $f$ is unchanged compared to before the update. Hence there can be at most one change per base in the collection, so $|\mathcal{B}|$ in total.
Since each update to a minimum weight base uses at most $O(\log(Wn))$ queries per update by Proposition˜1.1, we get a total of at most $O(|\mathcal{B}|^{2}\log(Wn))$ worst-case queries per update. Finally, note that $W\leq|\mathcal{B}|$ to obtain the results. ∎
B.2 Greedy Approximates Ideal Base Collection
@@
Now, we show that when we compute a greedy base collection $\mathcal{B}$ that is large enough, its relative loads $\ell^{\mathcal{B}}(\cdot)$ give good approximations to the ideal relative loads $\ell^{*}(\cdot)$. This generalizes the result for packing trees [Thorup07] to matroids, so the proof follows the approach used for the graph case. It utilizes a technique by Young [young95]. The idea is to consider a distribution of bases such that picking the bases for a base collection randomly from this distribution would result in the relative loads being ideal in expectation. Now, the number of violations, where the relative load of an element does not approximate its ideal load well, is analyzed. This is done while replacing the randomly picked trees one after the other by trees computed greedily. For the analysis, we consider a pessimistic estimator for the number of violations at every step of the process and show that its value cannot increase when a greedy tree is added.
Now, consider $\Pi_{\mathcal{M}}$, a distribution of bases of $\mathcal{M}$ constructed simultaneously with the assignment of the ideal relative loads. In the following, we generalize the proof of Lemma 13 from [Thorup07].
Lemma B.1.
@@
Let $\mathcal{B}^{*}$ be a base collection for $\mathcal{M}$ with $\frac{1}{\max_{e\in E}\ell^{\mathcal{B}^{*}}(e)}=\Phi_{\mathcal{M}}$. To pick a random base from $\Pi_{\mathcal{M}}$, pick a base $B$ from $\mathcal{B}^{*}$ uniformly at random. Then, pick an independent set $B^{\prime}$ randomly from the recursively defined distribution $\Pi_{\mathcal{M}|\overline{A_{0}}}$. We construct a base $B^{\prime\prime}:=(B\cap A_{0})\cup B^{\prime}$. First, we show that $B^{\prime\prime}$ is indeed a base for $\mathcal{M}$.
@@
The inequality follows from Lemma˜B.1 because every element $e\in A$ has ideal relative load $\ell^{*}(e)\leq d$. The next step is to show that the greedy base has a smaller quantity $q$ than for any base in $\Pi_{\mathcal{M}}$. We also have
@@
where the first inequality holds due to ˜B.3 and for the second inequality we note that, for any $B^{\prime}\in\Pi_{\mathcal{M}}$, we have that $B^{\prime}\cap A$ is a base for $\mathcal{M}|A$. Otherwise $B^{\prime}\cap A$ would not span $A$ and there would be an element $e\in A\setminus B^{\prime}$ such that $(B^{\prime}\cap A)+e\in\mathcal{I}$. By the exchange property we could add elements from $B$ to $(B^{\prime}\cap A)+e$ until we have another base for $\mathcal{M}$. Since the new base contains all elements from $B$ except for one element from $e^{\prime}\in E\setminus A$ with $\ell^{*}(e^{\prime})>d$ that is replaced with $e$ with $\ell^{*}(e)\leq d$ this contradicts that $B^{\prime}$ is a minimum weight base with respect to the ideal relative loads $\ell^{*}(\cdot)$. Therefore, for any base $B^{\prime\prime}$ of $\mathcal{M}|A$, specifically also for $B^{\prime}\cap A$, we have $c(B_{A})\leq c(B^{\prime\prime}\cap A)$ as $B_{A}$ is also a minimum weight base with respect to $(1+\varepsilon)^{L^{\mathcal{B}}(\cdot)}$ since the ordering between the elements stays the same if this value is considered instead of $L^{\mathcal{B}}(\cdot)$. ∎
@@
As a result, for a base $B_{A}$ as described in ˜B.7 we have

### v2

@@ 4.2 Base Covering
Let $E_{i}:=E\setminus\bigcup_{j<i}A_{j}$ and let $\mathcal{M}_{i}=\mathcal{M}|E_{i}$ for all recursion levels denoted by the indices $i\in I$. First, we note that $1/\min_{e}\ell^{*}(e)=\max_{i\in I}\Phi_{\mathcal{M}_{i}}$ over all recursion levels $i$. Let $j$ be the last iteration in which $\Phi_{\mathcal{M}_{i}}$ is maximized. Since the $\Phi$-values are non-decreasing, see Lemma˜4.2, this also has to be the last level of the recursion and therefore $E_{j+1}=E_{j}\setminus A_{j}=\emptyset$ and $\operatorname{rk}(E_{j})=\operatorname{rk}(A_{j})$. We get
@@
$|Y|=\sum_{i\in I}|A_{i}\cap Y|.$
Next, we want to bound $|Y|$ by $\sum_{i\in I}\Phi_{\mathcal{M}_{i}}\operatorname{rk}(A_{i}\cap Y)$ where it remains to show the inequality $|A_{i}\cap Y|\leq\Phi_{\mathcal{M}_{i}}\operatorname{rk}(A_{i}\cap Y)$ for any $i\in I$. Assume towards a contradiction that there is a level $k\in I$ for which the inequality does not hold. In the following let $\overline{A_{i}}:=E_{i}\setminus A_{i}$ for each $i\in I$. Now, consider the value
@@
since the matroid $\mathcal{M}_{i}$ contains $\Phi_{\mathcal{M}_{i}}$ disjoint bases. Hence, every base of $\mathcal{M}_{i}|(A_{i}\cap Y)$ contains at most $|A_{i}\cap Y|/\Phi_{\mathcal{M}_{i}}$ many elements. Next, we want to show
@@
5 Greedy Base Collection
First, let us define a greedy base collection as follows: let the weight of an element be the number of bases an element belongs to, and require that the base in the collection form successive minimum weight bases. This is a natural generalization of the notion for graphic matroids called ‘greedy tree-packing’ [^7] We opt for ‘collection’ instead of ‘packing’ to avoid any confusion with base packing as in Theorem 4.1., see, e.g., [Thorup07].
@@ 5 Greedy Base Collection
5.1 Dynamic Greedy Base Collection
@@
Given a greedy base collection $\mathcal{B}$, let there be an deletion or insertion $e$. First, consider the deletion of $e$. The element $e$ is part of at most $|\mathcal{B}|$ bases. To update our data structures, we need to delete it from every base it is part of. We do this in sequence and show that each such deletion leads to at most $|\mathcal{B}|$ updates to a base. Since each update to a minimum weight base uses at most $O(\log(n))$ queries per update by Proposition˜1.1, we get $O(|\mathcal{B}|^{2}\log(n))$ worst-case queries per update in total.
In the following, for any element $f\in E$ and $i\in\{1,\ldots,|\mathcal{B}|\}$, let $w_{i}(f)$ be the number of bases $B_{j}$ that contain $f$ with $j\leq i$.
It remains to show how to delete $e$ from a single base $B_{i}$ in $\mathcal{B}$. We do this by first updating $B_{i}$, then $B_{i+1}$ and so forth. We claim (by induction) that during this process, at any base $B_{j}$ (for $j\geq i_{e}$) we only need to make one update, i.e., increase the weight of one element $e^{\prime}$ and potentially replace it.
Let $B_{j}\in\mathcal{B}$ be the base that where we need to increase the weight of $e^{\prime}$. We claim that we get at most $|\mathcal{B}|-j$ changes to other bases. Note that this is non-trivial; at a first glance, the changes over bases could cascade. However, we note that if deleting $e^{\prime}$ from $B_{j}$ leads to adding $f$ to $B_{j}$, then $f$ gets an increased weight $w_{j}^{\text{new}}(f)=w_{j}^{\text{old}}(f)+1$. Now consider the next base $B_{k}$ with $k>j$ that contains $f$. We have two cases: either $f$ remains in the base $B_{k}$, even though its weight is increased. This means we have $w_{k}^{\text{new}}(f)=w_{k}^{\text{old}}(f)+1$ and all other elements still have the same weight. Alternatively, an element $g$ is replacing $f$ in $B_{k}$. This means that $w_{k}^{\text{new}}(g)=w_{k}^{\text{old}}(g)+1$ and $w_{k}^{\text{new}}(f)=w_{k}^{\text{old}}(f)+1-1=w_{k}^{\text{old}}(f)$. So for bases $B_{l}$ for $l>k$, we only need to adjust the weight of $g$. Hence, there can be at most one change per base in the collection, so $|\mathcal{B}|$ in total. The proof for an insertion of $e$ is analogous. ∎
5.2 Greedy Approximates Ideal Base Collection
@@
Now, we show that when we compute a greedy base collection $\mathcal{B}$ that is large enough, its relative loads $\ell^{\mathcal{B}}(\cdot)$ give good approximations to the ideal relative loads $\ell^{*}(\cdot)$. This generalizes the result for packing trees [Thorup07] to matroids, so the proof follows the approach used for the graph case. It utilizes a technique by Young [young95]. The idea is to consider a distribution of bases such that picking the bases for a base collection randomly from this distribution would result in the relative loads being ideal in expectation. Now, the number of violations, where the relative load of an element does not approximate its ideal load well, is analyzed. This is done while replacing the randomly picked bases one after the other by bases computed greedily. For the analysis, we consider a pessimistic estimator for the number of violations at every step of the process and show that its value cannot increase when a greedy base is added.
Now, we define $\Pi_{\mathcal{M}}$, a probability distribution of bases of $\mathcal{M}$. We construct $\Pi_{\mathcal{M}}$ simultaneously with the assignment of the ideal relative loads. See the process that defines ideal relative loads in Section˜1.1.1. On each (restricted) matroid $\mathcal{M}$ in this process, let $\mathcal{B}^{*}$ be a base collection for $\mathcal{M}$ with $\frac{1}{\max_{e\in E}\ell^{\mathcal{B}^{*}}(e)}=\Phi_{\mathcal{M}}$. To pick a random base from $\Pi_{\mathcal{M}}$, pick a base $B$ from $\mathcal{B}^{*}$ uniformly at random.
In the following, we generalize the proof of Lemma 13 from [Thorup07].
Lemma 5.1.
@@
First, pick a random base from $\Pi_{\mathcal{M}}$. Then, pick an independent set $B^{\prime}$ randomly from the recursively defined distribution $\Pi_{\mathcal{M}|\overline{A_{0}}}$. We construct a base $B^{\prime\prime}:=(B\cap A_{0})\cup B^{\prime}$. First, we show that $B^{\prime\prime}$ is indeed a base for $\mathcal{M}$.
@@
The inequality follows from Lemma˜5.1 because every element $e\in A$ has ideal relative load $\ell^{*}(e)\leq d$. The next step is to show that the greedy base has a smaller quantity $q$ than for any base in $\Pi_{\mathcal{M}}$. We also have
@@
where the first inequality holds due to ˜5.3 and for the second inequality we note that, for any $B^{\prime}\in\Pi_{\mathcal{M}}$, we have that $B^{\prime}\cap A$ is a base for $\mathcal{M}|A$. Otherwise $B^{\prime}\cap A$ would not span $A$ and there would be an element $e\in A\setminus B^{\prime}$ such that $(B^{\prime}\cap A)+e\in\mathcal{I}$. By the exchange property we could add elements from $B$ to $(B^{\prime}\cap A)+e$ until we have another base for $\mathcal{M}$. Since the new base contains all elements from $B$ except for one element from $e^{\prime}\in E\setminus A$ with $\ell^{*}(e^{\prime})>d$ that is replaced with $e$ with $\ell^{*}(e)\leq d$ this contradicts that $B^{\prime}$ is a minimum weight base with respect to the ideal relative loads $\ell^{*}(\cdot)$. Therefore, for any base $B^{\prime\prime}$ of $\mathcal{M}|A$, specifically also for $B^{\prime}\cap A$, we have $c(B_{A})\leq c(B^{\prime\prime}\cap A)$ as $B_{A}$ is also a minimum weight base with respect to $(1+\varepsilon)^{L^{\mathcal{B}}(\cdot)}$ since the ordering between the elements stays the same if this value is considered instead of $L^{\mathcal{B}}(\cdot)$. ∎
@@
As a result, for a base $B_{A}$ as described in ˜5.7 we have